OpenStudy (anonymous):

thitt=* the value of cos 1degree cos 2degree.........................cos 100degree is

OpenStudy (anonymous):

thitta=*\[\theta\]

OpenStudy (anonymous):

1 + 2cos x + 2 cos 2x + 2 cos 3x +....+ 2 cos nx =sin((n+(1/2))x) / sin(x/2)

OpenStudy (anonymous):

it is zero since cos90=0 and cos 90 appears in the series